OpenStudy (anonymous):

1 / x^-1 My mind just went blank? Can someone guide me?

OpenStudy (anonymous):

\[\frac{ 1 }{ x ^{-1} }\] ?

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

a negative exponent in the denominator is equivalent to a positive one in the numerator and vice versa. So \[\frac{ 1 }{ x ^{-1} } = x ^{1}\] and \[x ^{-1} = \frac{ 1 }{ x ^{1} }\]

OpenStudy (anonymous):

ohhh i see. thanks

OpenStudy (anonymous):

No problem, also, remember that anything raised to the power of 1 is just that thing. So, x^1 is just x